Calculate the pH of 1.0 x 10^-10 M Ca(OH)2
1
Expert's answer
2020-04-06T12:01:08-0400

"[OH^-]=2[Ca^{2+}]+[H^+]"

"K_w=[H +] [OH^-]=([OH^-]-2[Ca^{2+}])[OH^-]"

"[OH^-]^2-2[Ca^{2+}][OH^-]-K_w=0"

"[OH^-]^2-2*10^{-10}[OH^-]-10^{-14}=0"

"[OH]=1.001*10^{-7}"

"pOH=-lg[OH^-]=6.999"

"pH=K_w-pOH=14-6.999=7.001"

The answer is 7.001